# A Survey of Reinforcement Learning from Human Feedback

### v1

@@ 2. Unknown
[274] consider a similar learning scenario as [225] but with the additional objective to keep the number of queries of trajectory comparisons low, which is a combination of two competing objectives also studied in the bandit literature [67]. For this purpose, they suggest the Preference-based and Randomized Least-Squares Value Iteration (PR-LSVI) algorithm, which combines least-squares value iteration with a perturbed state-action-based reward model with Gaussian noise for regret minimization; a similar idea to CoLSTIM suggested for contextual dueling bandits [23]. More specifically, in each time step, the policy maximizing the value function of the perturbed state-action-based reward model and the policy maximizing the later in the previous time steps are “played”. By sampling trajectories of these two policies and computing their expected absolute reward difference (based on the perturbed state-action-based reward model) as a measure of uncertainty, preference feedback for these two trajectories is queried if the uncertainty exceeds a certain threshold. Moreover, they also suggest a posterior sampling counterpart of this algorithm, the Preference-based Thompson Sampling (PbTS) algorithm, and analyze it in terms of Bayesian quantities.
@@
[294]
Greedy policy for pessimistic expected value function estimation
Logistic link function, linear reward function for a state-pair feature embedding with some regularity assumptions on weights, known transition dynamics
High probability bound for the performance gap based on trajectory-based (and action-based) feedback
@@
[286]
Greedy policy for pessimistic expected value function estimation
General differentiable link function $\Phi$, general bounded reward function class $\mathcal{F}_{r}$ and general transition dynamic class $\mathcal{F}_{\mathbb{P}}$
High probability bound for the performance gap based on trajectory-based (and action-based) feedback
Transition dynamics:
@@
[161]
Value iteration based on pessimistic expected value function estimation
Dynamic discrete choice model, linear MDP, linear reward function for a state-pair feature embedding with some regularity assumptions on weights, known model class entailing the value and reward function of the dynamic discrete choice model
High probability bound for the performance gap based on action-based feedback
Linear model class:
@@
[145]
Maximum information gain for reward querying, value iteration based on pessimistic expected value function estimation for policy learning
Binary rewards for state-action pairs based on human response model with bounded noise, compliant and tabular MDP
High probability bound for the performance gap based on binary state-action reward feedback
Linear model class:
@@
[294] study the performance of a greedy policy trained from a data set consisting of trajectory pairs along with the observed preference that is assumed to be generated by means of a Bradley-Terry model with linear rewards. For this purpose, different results with respect to the MLE of the Bradley-Terry model for different feedback scenarios are derived that are quite of independent interest. In particular, they show concentration inequalities of the MLE for trajectory-based comparison feedback and additionally its asymptotic normality for action-based comparison feedback that also holds for $K$-wise comparisons. Based on these, it is shown that the greedy policy using the MLE in the case of action-based feedback might fail while using a pessimistic MLE leads to minimax-rates with respect to the performance gap [^7] The expected difference between the optimal value function and the value function of the used policy.. The latter is also shown to be true in the case of trajectory-based feedback. Technically, the pessimistic MLE is realized by taking the policy that has the largest pessimistic expected value function, i.e., the lowest realization of the value function within a hyperparameter-dependent confidence region around the MLE. Further results of independent interest are the inferred theoretical guarantees for maximum entropy inverse RL [295] and action-based inverse RL algorithms [193].
@@
The simple model assumptions underlying [294] were then replaced by more sophisticated assumptions in some subsequent work. The linear reward assumption has been replaced by more general reward function classes by [286] and [161]. In addition, [286] also consider more general unknown human feedback models and construct the confidence regions for the pessimistic approach directly from the log-likelihood function. The resulting approach, called FREEHAND, is analyzed in terms of its performance gap, for which some problem-dependent coefficients, the per-step, per-trajectory, and transition concentrability coefficient, are introduced. On the basis of a lower bound, it is shown that the per-trajectory concentrability coefficient should naturally appear in the bound on the performance gap. Moreover, the concentrability coefficient is shown to be upper bounded by the constant appearing in the special case of linear rewards considered by [294]. Finally, it is worth mentioning that both trajectory-based and action-based comparison feedback are considered.
@@
Assuming a dynamic discrete choice model [221] underlying the given data set of observed trajectories (without explicitly observed preferences), [161] suggest the Dynamic-Choice-Pessimistic-Policy-Optimization (DCPPO) algorithm. It first estimates the reward model using this assumption and then learns a policy in a (pessimistic) value iteration manner from the estimated reward model. In the case of a linear MDP and a known model class that entails both the value and the reward function of the dynamic discrete choice model, DCPPO is analyzed with respect to its performance gap. This is done for the case of a linear function model class as well as a subset of a reproducing kernel Hilbert space (RKHS) as the model class.
@@
Focusing on the estimation of the weight parameter in the Bradley-Terry model for the action-based feedback under label differential privacy conditions [72], [55] analyze two estimation procedures, MLE and stochastic gradient descent (SGD), under similar assumptions as in [294]. In both cases, the cost of ensuring label differential privacy is a multiplicative factor.
@@
Reward collapse, a term introduced by [243], describes the issue when rank-based training methods for LLMs lead to the same reward distribution regardless of the prompts used in the final training steps. The authors show that this occurs because the rank-based approach does not adequately account for prompt-related information. To address this problem, the authors propose a family of utility functions as well as an optimization method that successfully creates prompt-dependent reward distributions, effectively mitigating the collapse of rewards during training.
@@
[145] study the problem of optimal policy learning from critique feedback (see Section 3.2), i.e., binary rewards for state-action pairs, with as few queries to the human as possible. They assume an underlying ground-truth human feedback model that leads to a positive evaluation for a state-action pair if it exceeds a specific threshold evaluated at that pair. In addition, the learning process consists of two phases: First, exploring the environment in an unsupervised manner, and then querying user feedback in an active reward learning phase to learn the human feedback model. This learning process is again analyzed in two variants: Either the exploration phase was performed externally, and a data set consisting of trajectories is provided (offline), or this data set is actively collected itself (online). For both variants, an active learning algorithm is proposed that essentially selects query points (state-action pairs) that provide the most information gain given the points already designated to be queried. For the online variant, an exploration strategy based on optimistic least-squares value iteration [129] is also introduced for tabular or linear MDPs. In both variants, policy learning is carried out by a pessimistic value iteration with the empirical transitions and the estimated reward function, resulting in UCBVI-Planning (online) and LCBVI-Tabular-Offline (offline). Under the assumption of bounded noise [177] or low-noise assumption [147, 105], bounds on the performance gap of both algorithms are derived.
The question of the ideal experimental design for RLHF is addressed by [287], in particular, how to separate the process of data acquisition (e.g., trajectories to be evaluated) from the process of retrieving human feedback to avoid constantly involving humans in the training loop. Assuming linear rewards, the Bradley-Terry model and either a transition oracle (e.g., available for tabular or low-rank MDPs) or a linear MDP they suggest the expeRimental dEsiGn for queryIng huMan prEference (REGIME) algorithm that first samples exploratory trajectories indented to be as informative as possible for learning the reward via MLE and then applies a greedy policy based on the reward learned by the latter. They explicitly show that REGIME requires less human feedback to be queried in order to output an $\epsilon$-optimal policy at the end than the approach by [225].
@@
There have been some theoretical analyses regarding the question in how far, or if at all, preference-based feedback in the form of trajectory comparisons is more s more suitable compared to numerical feedback. [127] suggest a human rating model for this purpose in the numerical feedback case and analyze the LCB algorithm [130] in order to compare it with the pessimistic MLE [294]. It is shown that under specific assumptions, LCB has a constant performance gap, while the preference-based pessimistic MLE under similar assumptions has a similar bound as in Table 6.
[262] provide reduction-based algorithms that can directly utilize state-of-the-art results in reward-based RL for RLHF with utility-based and general state-action and trajectory-based comparison feedback. They show, in general, how theoretical results of the underlying standard RL algorithm can be translated to theoretical results for the resulting preference-based RL algorithm. For some special cases, such as MDPs with finite Eluder dimension and utility-based preference feedback, the theoretical guarantees are explicitly derived using state-of-the-art RL algorithms that are qualitatively similar to explicit preference-based RL algorithms.
@@ 7 Applications and Benchmarks
The field of RLHF has advanced significantly in the last few years, with increasing interest driven by prominent applications. First and foremost are applications to large language models, exemplified by ChatGPT [197]. This section starts by providing a sample of such applications, showcasing how this technology is being utilized in fields as varied as robotics, language processing, image generation, and more. We will also delve into libraries that provide foundational support for RLHF research, enabling researchers and practitioners to experiment with and refine a range of approaches. We then explore a spectrum of benchmarks that have been developed to standardize and simplify the evaluation of new approaches, offering insights into their performance in different settings. Finally, and closely related to those benchmarks, we will discuss common evaluation practices.
@@ 7.1 Applications
RLHF finds applications across various domains, showcasing its versatility in addressing complex and nuanced tasks. The most prominent application is ChatGPT [197], which is an example of an application in the domain of language models. Beyond that, however, applications extend across diverse domains such as control tasks, generative models, and recommender systems. This section provides an overview of notable works applying RLHF in different areas.
@@
There is a long history of using control environments as benchmark tasks for RL. In addition to the breadth of available environments, control applications are of particular interest because tasks are often hard to specify. [58] demonstrated the effectiveness RLHF in games as well as simulated continuous control tasks, matching the performance of RL agents trained on ground-truth rewards with a fraction of the feedback. Extending to robotics, [68] trained a reward model for diverse tasks with a single robot, achieving human-like behavior. [149] applied RLHF for precise robot-to-human handovers. Similarly, [2] used RLHF in the Playhouse simulator, a platform for sensorimotor task training, and [182] showcase an application in the context of the MineRL Basalt competition for Minecraft tasks.
@@
Generative models, i.e., models that generate new data instead of just predicting labels, can be framed as an RL setting in which a policy assembles the output through its actions. In the context of language models, this means that the language model is interpreted as a policy with tokens as actions. Using this reframing, we can use RLHF approaches to fine-tune generative models to produce preferred outputs. ChatGPT [197] and GPT-4 [198] are prime examples of language models fine-tuned using RLHF. These applications build on earlier work, such as by [200], [296] and [96]. This method extends to text summarization [87, 88, 244], dialogue summarization [50], and question answering [191]. In image generation, [155] and [279] demonstrate the use of reward modeling for text-to-image tasks, while [204] and [137] explore RLHF applications in broader computer vision tasks.
@@
In the context of recommender systems, [282] have shown the potential of RLHF in optimizing for long-term engagement. Although it is, in principle, possible to algorithmically evaluate policies in this domain, these rewards are sparse. To combat this, [282] use RLHF to distill sparse, global feedback into a dense reward model.

### v2

@@ 2. Unknown
Wu & Sun (2024) consider a similar learning scenario as Saha et al. (2023) but with the additional objective to keep the number of queries of trajectory comparisons low, which is a combination of two competing objectives also studied in the bandit literature (Degenne et al., 2019). For this purpose, they suggest the Preference-based and Randomized Least-Squares Value Iteration (PR-LSVI) algorithm, which combines least-squares value iteration with a perturbed state-action-based reward model with Gaussian noise for regret minimization; a similar idea to CoLSTIM suggested for contextual dueling bandits (Bengs et al., 2022). More specifically, in each time step, the policy maximizing the value function of the perturbed state-action-based reward model and the policy maximizing the later in the previous time steps are “played”. By sampling trajectories of these two policies and computing their expected absolute reward difference (based on the perturbed state-action-based reward model) as a measure of uncertainty, preference feedback for these two trajectories is queried if the uncertainty exceeds a certain threshold. Moreover, they also suggest a posterior sampling counterpart of this algorithm, the Preference-based Thompson Sampling (PbTS) algorithm, and analyze it in terms of Bayesian quantities.
Recently, Chakraborty et al. (2024) proposed a bilevel optimization problem that generalizes the standard optimization problem for RLHF with trajectory feedback and the negative log-likelihood as a loss function (i.e., (2)). This problem, which they call PARL (Policy Alignment in Reinforcement Learning), is characterized by explicitly taking into account the dependence on the data-collecting process at one level for the optimal policy parameters at the other level. For this problem, A-PARL is proposed, which is shown to have an $O(1/T)$ convergence rate under specific assumptions, where $T$ is the number of iterations.
Finally, for the LLM training scenario, there are two perspectives on the problem (MDP vs. contextual dueling bandits) as mentioned in Section 6. Accordingly, work from the field of contextual preference-based bandits (Dudík et al., 2015; Saha, 2021; Saha & Krishnamurthy, 2022; Bengs et al., 2022; Sekhari et al., 2023) can also be viewed as theoretical contributions for the RLHF setting for the LLM application. This view is advocated in particular by Xiong et al. (2024), who view the LLM fine-tuning task as a reverse-KL regularized contextual bandit problem. Typically the context is chosen externally, but Mehta et al. (2023) consider the learning variant in which the learning agent chooses the context as well. This variant is referred to as active contextual dueling bandits.
@@
(Zhu et al., 2023a)
Greedy policy for pessimistic expected value function estimation Logistic link function, linear reward function for a state-pair feature embedding with some regularity assumptions on weights, known transition dynamics High probability bound for the performance gap based on trajectory-based (and action-based) feedback
@@
(Zhan et al., 2024a)
Greedy policy for pessimistic expected value function estimation General differentiable link function $\Phi$, general bounded reward function class $\mathcal{F}_{r}$ and general transition dynamic class $\mathcal{F}_{\mathbb{P}}$ High probability bound for the performance gap based on trajectory-based (and action-based) feedback Transition dynamics:
@@
(Li et al., 2023)
Value iteration based on pessimistic expected value function estimation Dynamic discrete choice model, linear MDP, linear reward function for a state-pair feature embedding with some regularity assumptions on weights, known model class entailing the value and reward function of the dynamic discrete choice model High probability bound for the performance gap based on action-based feedback Linear model class:
@@
(Kong & Yang, 2022)
Maximum information gain for reward querying, value iteration based on pessimistic expected value function estimation for policy learning Binary rewards for state-action pairs based on human response model with bounded noise, compliant and tabular MDP High probability bound for the performance gap based on binary state-action reward feedback Linear model class:
@@
Zhu et al. (2023a) study the performance of a greedy policy trained from a data set consisting of trajectory pairs along with the observed preference that is assumed to be generated by means of a Bradley-Terry model with linear rewards. For this purpose, different results with respect to the MLE of the Bradley-Terry model for different feedback scenarios are derived that are quite of independent interest. In particular, they show concentration inequalities of the MLE for trajectory-based comparison feedback and additionally its asymptotic normality for action-based comparison feedback that also holds for $K$-wise comparisons. Based on these, it is shown that the greedy policy using the MLE in the case of action-based feedback might fail while using a pessimistic MLE leads to minimax-rates with respect to the performance gap [^8] The expected difference between the optimal value function and the value function of the used policy.. The latter is also shown to be true in the case of trajectory-based feedback. Technically, the pessimistic MLE is realized by taking the policy that has the largest pessimistic expected value function, i.e., the lowest realization of the value function within a hyperparameter-dependent confidence region around the MLE. Further results of independent interest are the inferred theoretical guarantees for maximum entropy inverse RL (Ziebart et al., 2008) and action-based inverse RL algorithms (Neu & Szepesvári, 2009).
@@
The simple model assumptions underlying (Zhu et al., 2023a) were then replaced by more sophisticated assumptions in some subsequent work. The linear reward assumption has been replaced by more general reward function classes by Zhan et al. (2024a) and Li et al. (2023). In addition, Zhan et al. (2024a) also consider more general unknown human feedback models and construct the confidence regions for the pessimistic approach directly from the log-likelihood function. The resulting approach, called FREEHAND, is analyzed in terms of its performance gap, for which some problem-dependent coefficients, the per-step, per-trajectory, and transition concentrability coefficient, are introduced. On the basis of a lower bound, it is shown that the per-trajectory concentrability coefficient should naturally appear in the bound on the performance gap. Moreover, the concentrability coefficient is shown to be upper bounded by the constant appearing in the special case of linear rewards considered by Zhu et al. (2023a). Finally, it is worth mentioning that both trajectory-based and action-based comparison feedback are considered.
In follow-up work, Zhu et al. (2024) found overfitting as well as overoptimization issues of the MLE in the Boltzmann model for pairwise comparison feedback. This can arise in particular if the observations of labels are strongly unbalanced and thus the utilities can become infinite. To overcome this problem, they propose the Iterative Data Smoothing (IDS) algorithm, which implicitly weights observed labels appropriately by their frequency and their current likelihood. Note that these issues do not contradict the results shown by Zhu et al. (2023a) as these are based on the assumption of bounded utilities (or rewards).
@@
Assuming a dynamic discrete choice model (Rust, 1987) underlying the given data set of observed trajectories (without explicitly observed preferences), Li et al. (2023) suggest the Dynamic-Choice-Pessimistic-Policy-Optimization (DCPPO) algorithm. It first estimates the reward model using this assumption and then learns a policy in a (pessimistic) value iteration manner from the estimated reward model. In the case of a linear MDP and a known model class that entails both the value and the reward function of the dynamic discrete choice model, DCPPO is analyzed with respect to its performance gap. This is done for the case of a linear function model class as well as a subset of a reproducing kernel Hilbert space (RKHS) as the model class.
@@
Focusing on the estimation of the weight parameter in the Bradley-Terry model for the action-based feedback under label differential privacy conditions (Dwork, 2008), Chowdhury & Zhou (2023) analyze two estimation procedures, MLE and stochastic gradient descent (SGD), under similar assumptions as in Zhu et al. (2023a). In both cases, the cost of ensuring label differential privacy is a multiplicative factor.
@@
Reward collapse, a term introduced by Song et al. (2023), describes the issue when rank-based training methods for LLMs lead to the same reward distribution regardless of the prompts used in the final training steps. The authors show that this occurs because the rank-based approach does not adequately account for prompt-related information. To address this problem, the authors propose a family of utility functions as well as an optimization method that successfully creates prompt-dependent reward distributions, effectively mitigating the collapse of rewards during training.
@@
Kong & Yang (2022) study the problem of optimal policy learning from critique feedback (see Section 3.2), i.e., binary rewards for state-action pairs, with as few queries to the human as possible. They assume an underlying ground-truth human feedback model that leads to a positive evaluation for a state-action pair if it exceeds a specific threshold evaluated at that pair. In addition, the learning process consists of two phases: First, exploring the environment in an unsupervised manner, and then querying user feedback in an active reward learning phase to learn the human feedback model. This learning process is again analyzed in two variants: Either the exploration phase was performed externally, and a data set consisting of trajectories is provided (offline), or this data set is actively collected itself (online). For both variants, an active learning algorithm is proposed that essentially selects query points (state-action pairs) that provide the most information gain given the points already designated to be queried. For the online variant, an exploration strategy based on optimistic least-squares value iteration (Jin et al., 2020) is also introduced for tabular or linear MDPs. In both variants, policy learning is carried out by a pessimistic value iteration with the empirical transitions and the estimated reward function, resulting in UCBVI-Planning (online) and LCBVI-Tabular-Offline (offline). Under the assumption of bounded noise (Massart & Nédélec, 2006) or low-noise assumption (Korba et al., 2017; Haddenhorst et al., 2021), bounds on the performance gap of both algorithms are derived.
The question of the ideal experimental design for RLHF is addressed by Zhan et al. (2024b), in particular, how to separate the process of data acquisition (e.g., trajectories to be evaluated) from the process of retrieving human feedback to avoid constantly involving humans in the training loop. Assuming linear rewards, the Bradley-Terry model and either a transition oracle (e.g., available for tabular or low-rank MDPs) or a linear MDP they suggest the expeRimental dEsiGn for queryIng huMan prEference (REGIME) algorithm that first samples exploratory trajectories indented to be as informative as possible for learning the reward via MLE and then applies a greedy policy based on the reward learned by the latter. They explicitly show that REGIME requires less human feedback to be queried in order to output an $\epsilon$-optimal policy at the end than the approach by Saha et al. (2023).
@@
There have been some theoretical analyses regarding the question in how far, or if at all, preference-based feedback in the form of trajectory comparisons is more suitable compared to numerical feedback. Ji et al. (2023c) suggest a human rating model for this purpose in the numerical feedback case and analyze the LCB algorithm (Jin et al., 2021) in order to compare it with the pessimistic MLE (Zhu et al., 2023a). It is shown that under specific assumptions, LCB has a constant performance gap, while the preference-based pessimistic MLE under similar assumptions has a similar bound as in Table 7.
Wang et al. (2023b) provide reduction-based algorithms that can directly utilize state-of-the-art results in reward-based RL for RLHF with utility-based and general state-action and trajectory-based comparison feedback. They show, in general, how theoretical results of the underlying standard RL algorithm can be translated to theoretical results for the resulting preference-based RL algorithm. For some special cases, such as MDPs with finite Eluder dimension and utility-based preference feedback, the theoretical guarantees are explicitly derived using state-of-the-art RL algorithms that are qualitatively similar to explicit preference-based RL algorithms.
7.3 Nash Learning from Human Feedback
The majority of theoretical works use the modeling of (pairwise comparison) feedback by means of a link function (see Section 5.1.4). Even if this often leads to simpler derivations, this modeling has the decisive disadvantage that it imposes a transitivity of the human feedback that does not necessarily prevail in reality. In other words, it is quite possible that preference cycles can occur. For this reason, there is a new direction in theoretical work that dispenses with parametric modeling of the preference probability similar to Chen et al. (2022) but uses it to formulate a new learning objective. Specifically, the problem is considered from a game theory perspective, where two policies each propose a trajectory that should be highly preferred by the human user. Thus, the goal is to find a policy that suggests trajectories that are preferred to the trajectories of any other policy, i.e., a Nash equilibrium or a von Neumann winner.
This learning variant was first considered by Wang et al. (2023b), who showed that the problem can be reduced to finding restricted Nash equilibria in a multi-agent RL problem (based on numerical rewards). For special situations of the latter problem, wrapper algorithms are proposed that have been shown to find the von Neumann winner with high probability. The learning problem was recently taken up and analyzed by Munos et al. (2023) and Ye et al. (2024) in a KL-regularization variant. While the former considers the online learning setting assuming a known preference model, the latter considers both online as well as offline learning settings and the preference model belonging to a finite function class.
@@ 8 Applications and Benchmarks
The field of RLHF has advanced significantly in the last few years, with increasing interest driven by prominent applications. First and foremost are applications to large language models, exemplified by ChatGPT (OpenAI, 2022). This section starts by providing a sample of such applications, showcasing how this technology is being utilized in fields as varied as robotics, language processing, image generation, and more. We will also delve into libraries that provide foundational support for RLHF research, enabling researchers and practitioners to experiment with and refine a range of approaches. We then explore a spectrum of benchmarks that have been developed to standardize and simplify the evaluation of new approaches, offering insights into their performance in different settings. Finally, and closely related to those benchmarks, we will discuss common evaluation practices.
@@ 8.1 Applications
RLHF finds applications across various domains, showcasing its versatility in addressing complex and nuanced tasks. The most prominent application is ChatGPT (OpenAI, 2022), which is an example of an application in the domain of language models. Beyond that, however, applications extend across diverse domains such as control tasks, generative models, and recommender systems. This section provides an overview of notable works applying RLHF in different areas.
@@
There is a long history of using control environments as benchmark tasks for RL. In addition to the breadth of available environments, control applications are of particular interest because tasks are often hard to specify. Christiano et al. (2017) demonstrated the effectiveness of RLHF in games as well as simulated continuous control tasks, matching the performance of RL agents trained on ground-truth rewards with a fraction of the feedback. Extending to robotics, Ding et al. (2023) trained a reward model for diverse tasks with a single robot, achieving human-like behavior. Kupcsik et al. (2018) applied RLHF for precise robot-to-human handovers. Similarly, Abramson et al. (2022) used RLHF in the Playhouse simulator, a platform for sensorimotor task training, and Milani et al. (2022) showcase an application in the context of the MineRL Basalt competition for Minecraft tasks. Recently, Dong et al. (2024) use RLHF to guide a diffusion-based planning model.
@@
Generative models, i.e., models that generate new data instead of just predicting labels, can be framed as an RL setting in which a policy assembles the output through its actions. In the context of language models, this means that the language model is interpreted as a policy with tokens as actions. Using this reframing, we can use RLHF approaches to fine-tune generative models to produce preferred outputs. ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023) are prime examples of language models fine-tuned using RLHF. These applications build on earlier work, such as by Ouyang et al. (2022), Ziegler et al. (2020) and Glaese et al. (2022). This method extends to text summarization (Gao et al., 2018; 2020; Stiennon et al., 2020), dialogue summarization (Chen et al., 2023), and question answering (Nakano et al., 2022). In image generation, Lee et al. (2023) and Xu et al. (2023) demonstrate the use of reward modeling for text-to-image tasks, while Pinto et al. (2023) and Kazemi et al. (2020) explore RLHF applications in broader computer vision tasks. Interestingly, in the context of LLMs, reward learning has also been expressed as density estimation (Dumoulin et al., 2024) instead of the supervised approach described in Section 5.
@@
In the context of recommender systems, Xue et al. (2023b) have shown the potential of RLHF in optimizing for long-term engagement. Although it is, in principle, possible to algorithmically evaluate policies in this domain, these rewards are sparse. To combat this, Xue et al. (2023b) use RLHF to distill sparse, global feedback into a dense reward model.